OpenStudy (anonymous):

The square root of a product is equal to the product of the square roots of each factor. True False

OpenStudy (king):

yes.... \[\sqrt{ab}=\sqrt{a}\times \sqrt{b}\]

OpenStudy (king):

do u understand gsj?

OpenStudy (anonymous):

Thanks, now I understand.

OpenStudy (king):

yw!

OpenStudy (zarkon):

you need \(a,b\ge0\)

Directrix (directrix):

Over the set of Real numbers, the square root of a product is equal to the product of the square roots of each factor.

OpenStudy (zarkon):

\[1=\sqrt{1}=\sqrt{(-1)(-1)}\neq\sqrt{-1}\sqrt{-1}=i\cdot i=-1\]

OpenStudy (anonymous):

Thanks to all of you. True is the right answer.

OpenStudy (zarkon):

unfortunately, true is not the correct answer...unless you restrict what you put under the square root...as is evident by my last post.